Olive wants to invest dollar-sign Baseline 15 Number , 000 with the goal of earning 10 % interest. She wants to put some of the money in an investment that has been earning 7 % a year and the rest of the money in an investment that typically earns 11 % a year. How much money should be invested at 11 % interest for Olive to obtain her goal of an overall 10 % interest?
Answer by stanbon(75887)

Olive wants to invest 15,000 with the goal of earning 10 % interest. She wants to put some of the money in an investment that has been earning 7 % a year and the rest of the money in an investment that typically earns 11 % a year. How much money should be invested at 11 % interest for Olive to obtain her goal of an overall 10 % interest?
S + E = 15000 dollars
0.07S + 0.11E = 0.10*15000
------
7S + 7E = 7*15000
7S + 11E = 10*15000
--------------------------
Subtract and solve for "E"::
4E = 3*15000
E = $11250 (amt. to invest at 11%)
S = 15000-11250 = $2750 (amt. to invest at 7%)
-----------------
Cheers,
Stan H.
-------------